### v1

@@ Abstract
Spatial heterogeneity widely exists in many applications, such as in ocean science, where the temperature-salinity (T-S) relationship in thermohaline circulation varies across different geographical locations and depths. While spatial regression models are powerful tools for this purpose, they often face challenges in simultaneously estimating spatial parameters, detecting heterogeneity boundaries, and performing adaptive modeling, especially in complex systems. This paper proposes a Fused Lasso regression model with an Adaptive minimum spanning Tree (FLAT) to address these challenges in a unified framework. Specifically, FLAT constructs an adaptive minimum spanning tree guided by both spatial proximity and coefficient dissimilarity, and incorporates a spatial heterogeneity penalty to capture the underlying structure. A subsequent spatial clustering algorithm then identifies discrete heterogeneity boundaries, such as oceanic thermohaline fronts. Numerical simulations confirm that FLAT significantly outperforms classic spatial regression models in both coefficient estimation and heterogeneity detection. An empirical analysis with Atlantic Ocean data further demonstrates FLAT’s capability to elucidate region-specific thermohaline compensation mechanisms and to detect surfaces with inverse T-S relationships. These findings advance the mechanistic understanding of T-S compensation dynamics in the Antarctic Intermediate Water region.
@@
Spatial heterogeneity , Fused regularization , Minimum spanning tree , Thermohaline circulation , Temperature-salinity relationship
11footnotetext:   All authors equally contribute.22footnotetext: * Corresponding author. Email: zxing@nd.edu
@@ 1 Introduction
Spatial heterogeneity is a fundamental feature of geophysical processes in the modern earth system, profoundly affecting phenomena from urban heat islands to the oceanic thermohaline circulation (THC) (Anselin, 2013; Talley, 2011). The THC, a global conveyor belt driven by density gradients from temperature and salinity variations, plays a pivotal role in regulating Earth’s climate (Clark et al., 2002). Its meridional overturning component transports vast amounts of heat poleward and is intrinsically linked to critical diapycnal mixing processes (Talley, 2013). A key challenge in oceanography is that this system exhibits a dual nature of spatial non-stationarity: it features both smooth, continuous variations within water masses and sharp, discrete transitions at their boundaries (Fotheringham et al., 2017; Ferrari et al., 2001). Accurately capturing this duality is crucial for understanding oceanic dynamics, yet it poses a significant challenge for existing analytical methods.
On one hand, spatial systems exhibit continuous heterogeneity, where spatial variables and their relationships vary smoothly across geographic space (Bailey et al., 1995). This phenomenon, also known as spatial continuous non-stationarity (Fotheringham et al., 2017), is evident in numerous environmental contexts. For instance, in climatology, temperature tends to decline and precipitation increases with elevation due to reduced atmospheric pressure and diminished heat absorption capacity. Similarly, in terrestrial hydrology, soil moisture-temperature relationships often vary continuously across space (Zhang et al., 2024). However, focusing solely on this smooth aspect of reality provides an incomplete picture, as many systems are also structured by abrupt, structural divisions. On the other hand, these abrupt divisions manifest as spatially discrete heterogeneity, referring to sharp transitions or categorical differences in variable relationships across space (Ferrari et al., 2001; Anselin, 2010). Such discontinuities are commonly observed in dynamic systems, from sudden shifts in rainfall regimes along mountain ridges (Luo et al., 2021) to sharp thermohaline interfaces marked by steep temperature gradients in the ocean (Li and Yang, 2021). Therefore, The fundamental scientific challenge lies in modeling both continuous and discrete patterns that coexist within the same geographic space.
@@
To model such heterogeneity, several spatial regression models and their extensions have been developed (LeSage and Pace, 2009; Ward and Gleditsch, 2018). Classic approaches include the spatial lag model (Lam and Souza, 2020), the spatial Durbin model (Anselin, 2013), and geographically weighted regression (GWR), which captures local spatial dependencies (Brunsdon et al., 1996). However, these models are primarily designed to estimate spatially varying coefficients, thereby capturing only the continuous aspect of heterogeneity. They generally do not aim to detect abrupt changes or spatial boundaries in the coefficient surfaces that are of central importance in fields like oceanography, such as the inverse temperature-salinity surface in thermohaline circulation. To address such cases, some empirical studies employ two-step procedures that involve space partitioning (Chen et al., 2024; Duque et al., 2011) and clustering-based segmentation techniques (Raudenbush, 2002; Van Dusen and Nissen, 2019). However, classic spatial models (Ward and Gleditsch, 2018) often fail to capture sharp transitions stably at regional boundaries. Since real-world geographic spaces often exhibit both continuous and discrete forms of spatial heterogeneity, it is crucial to develop adaptive modeling methods with capability to capture local smoothness while also identifying abrupt spatial boundaries of coefficients.
We propose a novel spatial regression framework, the Fused Lasso with an Adaptive minimum spanning Tree (FLAT), which addresses the challenges by introducing a structured adaptive regularization that captures both spatial heterogeneity in coefficients and sparsity in high-dimensional features. Specifically, we construct fused penalization terms guided by minimum spanning tree structures that incorporate both spatial proximity and coefficient similarity, and we incorporate adaptive weights to enhance estimation accuracy and computational efficiency. In addition, we include extra penalization terms to achieve simultaneous variable selection. A density-based spatial clustering method can be conducted to further detect the non-convex spatial regions with heterogeneity in applications, such as the detection of inverse temperature-salinity surface in the ocean. Comprehensive simulations and a real-world application to thermohaline circulation data demonstrate that FLAT yields accurate estimation, robust feature selection, and can provide additional unique insights in empirical analysis with interpretable results, which is applicable for data-driven discovery in geophysical sciences and other interdisciplinary applications.
@@
The rest of the article is organized as follows. Section 2 introduces the methodology of the fused Lasso regression with adaptive minimum spanning tree. Simulation studies are provided in Section 3, and a real application of our method on thermohaline circulation is shown in Section 4. Section 5 concludes this article with a discussion. Technical details and additional details are included in the Appendix.
@@ 2 Fused Lasso Regression With Adaptive Minimum Spanning Tree
To characterize both the complex relationships and heterogeneity in spatial data, we denote the spatial data at the $i$-th spatial location $s_{i}$ as $\{\bm{x}(s_{i}),y(s_{i})\}$ for $i\in[n]$, where $\bm{x}(s_{i})=\left[x_{1}(s_{i}),\cdots,x_{p}(s_{i})\right]^{\top}$ is the $p$-dimensional vector, $y(s_{i})$ is the corresponding response and $[n]=\{1,\cdots,n\}$. Then we consider a spatially varying coefficient regression model as below:
$y(s_{i})=\bm{\beta}^{\top}(s_{i})\mathbf{x}(s_{i})+\epsilon(s_{i}),$ (1)
[…]
@@
We firstly generate the minimum spanning tree $\mathbb{E}_{0}$ based on spatial distances and initialize the parameter estimators $\hat{\bm{\beta}}^{0}(s_{i}),i\in[n]$ by optimizing the following target function
@@
$\mathop{\arg\min}\limits_{\bm{\beta}}\frac{1}{n}\sum\limits_{i=1}^{n}\left(y(s_{i})-\bm{\beta}^{\top}(s_{i})\mathbf{x}(s_{i})\right)^{2}+\lambda\sum\limits_{k=1}^{p}\sum\limits_{(s_{i},s_{j})\in\mathbb{E}_{0}}|\beta_{k}(s_{i})-\beta_{k}(s_{j})|.$ (2)
Then we generate the full connection graph among spatial locations based on the $L_{2}$ distance between coefficients as
$\tilde{d}(s_{i},s_{j})=\|\hat{\bm{\beta}}^{0}(s_{i})-\hat{\bm{\beta}}^{0}(s_{j})\|_{2},$ (3)
and obtain the spatial linkage structure $\mathbb{E}$ via Prim’s algorithm based on the estimated distance matrix.
To introduce discrete spatial heterogeneity in the spatial regression model for continuous spatial heterogeneity, we construct the connectivity matrix $\mathbf{H}\in\mathbb{R}^{(n-1)\times n}$ from $\mathbb{E}$, which represents the minimum spanning tree formed by $(n-1)$ edges and $n$ vertices. The elements $\mathbf{H}_{l,i}$ of $\mathbf{H}$ are defined based on the connection status of vertices $i(1,...,n)$ along the edges $l(1,...,n-1)$. If the $l$-th edge $(s_{i},s_{j})\in\mathbb{E}$, then $\mathbf{H}_{l,i}=1$, $\mathbf{H}_{l,j}=-1$, and $\mathbf{H}_{l,v}=0$ for $v\neq i,j$. Thus, we have
$\|\mathbf{H}\bm{\beta}_{k}\|_{1}=\sum_{(s_{i},s_{j})\in\mathbb{E}}|\bm{\beta}_{k}(s_{i})-\bm{\beta}_{k}(s_{j})|,$ (4)
where $\bm{\beta}_{k}=(\beta_{k}(s_{1}),...,\beta_{k}(s_{n}))^{\top},\quad k=1,...,p$ represent the coefficient values of the $k$-th explanatory variable at each observation point. Based on equation (4), $\|\mathbf{H}\bm{\beta}_{k}\|_{1}$ represents the heterogeneity penalty based on the adaptive minimum spanning tree for the $k$-th explanatory variable.
To further improve the finite sample performance of both estimations and detection of discrete spatial heterogeneity, we choose the adaptive weight between any two spatial $(s_{i},s_{j})$ as $\pi_{(s_{i},s_{j})}=1/\tilde{d}(s_{i},s_{j})^{\gamma}$ for all $(s_{i},s_{j})\in\mathbb{E}$ and 00 otherwise. We follow the recommendation of Zou (2006) and set $\gamma=1$ , where $\gamma$ is a constant adjusting the adaptiveness of the penalty. Intuitively, spatial locations with small distances in the initial coefficients, denoted by $\tilde{d}(s_{i},s_{j})$, are more likely to belong to the same region or share similar underlying processes. In this context, larger adaptive weights impose stronger penalization on discrepancies between nearby coefficients, thereby encouraging local homogeneity and smoothing. This leads to estimated coefficients that are more likely to align in regions of weak heterogeneity, effectively capturing both smooth transitions and abrupt changes. Based on this intuition, the adaptive penalization on spatial heterogeneity is further formulated as:
where $\mathbb{E}$ represents the edge set derived from the minimum spanning tree based on the correlation similarity matrix, and $\mathcal{N}_{i}$ denotes the neighboring nodes of node $i$ in the minimum spanning tree. Ultimately, the penalty for spatial heterogeneity can be rewritten as:
$P_{\lambda_{1}}(\bm{\beta})=\lambda_{1}\sum_{k=1}^{p}\|\bm{\pi}\mathbf{H}\bm{\beta}_{k}\|_{1},$ (6)
where the weight matrix $\bm{\pi}\in\mathbb{R}^{(n-1)\times(n-1)}$ represents the weight of the edge $(s_{i},s_{j})$.
Regarding the sparsity of high-dimensional spatial features, we conduct variable selection via additional Lasso penalty to shrink coefficients to zero and to decrease model complexity. The objective function of FLAT is
$L_{\text{FLAT}}(\bm{\beta})=\frac{1}{n}\sum_{i=1}^{n}\left(y(s_{i})-\bm{\beta}^{\top}(s_{i})\mathbf{x}(s_{i})\right)^{2}+\lambda_{1}\sum_{k=1}^{p}\|\bm{\pi}\mathbf{H}\bm{\beta}_{k}\|_{1}+\lambda_{2}\sum_{k=1}^{p}\|\bm{\beta}_{k}\|_{1},$ (7)
where $\lambda_{1}$ and $\lambda_{2}$ are tuning parameters controlling the strength of penalization to capture the discrete spatial heterogeneity and the sparsity in spatial features. The adaptive weights $\bm{\pi}$ further enhance the detection of spatial heterogeneity while reducing the estimation error for the spatial locations with homogeneity.
For simplicity, we use matrix notation and rewrite the objective function of FLAT (7) as
$L_{\text{FLAT}}(\bm{\beta})=\frac{1}{2}\|\mathbf{y}-\mathbf{X}\bm{\beta}\|_{2}^{2}+\lambda_{1}\sum_{k=1}^{p}\|\bm{\pi}\mathbf{H}\bm{\beta}_{k}\|_{1}+\lambda_{2}\sum_{k=1}^{p}\|\bm{\beta}_{k}\|_{1},$ (8)
where $\mathbf{y}=[y(s_{1}),...,y(s_{n})]^{\top}\in\mathbb{R}^{n}$ represents the vector of response values for $n$ locations. $\mathbf{X}\triangleq[\text{diag}(\mathbf{x}_{1}),...,\text{diag}(\mathbf{x}_{p})]\in\mathbb{R}^{n\times np}$, where $\mathbf{x}_{k}$ represents the $k$-th covariate vector of length $n$, and the observation vector is $\mathbf{x}_{k}\triangleq(x_{k}(s_{1}),...,x_{k}(s_{n}))^{\top}\in\mathbb{R}^{n}$ is the corresponding design matrix with spatial feature. We denote $\bm{\beta}=[\bm{\beta}^{\top}(s_{1}),...,\bm{\beta}^{\top}(s_{n})]^{\top}\in\mathbb{R}^{np}$, where $\bm{\beta}(s_{i})\triangleq[\beta_{1}(s_{i}),...,\beta_{p}(s_{i})]^{\top}\in\mathbb{R}^{p}$.
To further simplify the notation, we define the matrix $\mathbf{\tilde{H}}$ and reparameterize coefficient matrix accordingly
$\mathbf{\tilde{H}}\triangleq\begin{bmatrix}\pi\mathbf{H}\\ \frac{\lambda_{2}}{\lambda_{1}}\mathbf{1}^{\top}\end{bmatrix}\in\mathbb{R}^{n\times n},\leavevmode\nobreak\ \leavevmode\nobreak\ \bm{\theta}_{k}=\begin{bmatrix}\pi\mathbf{H}\\ \frac{\lambda_{2}}{\lambda_{1}}\mathbf{1}^{\top}\end{bmatrix}\bm{\beta}_{k}=\mathbf{\tilde{H}}\bm{\beta}_{k},$ (9)
where $\bm{\theta}_{k}\triangleq[\theta_{k}(s_{1}),...,\theta_{k}(s_{n})]^{\top}\in\mathbb{R}^{n}$. Since $\mathbf{\tilde{H}}$ is invertible, we reparameterize $\bm{\beta}$ as equation (9) and express $\bm{\beta}_{k}$ as $\bm{\beta}_{k}=\mathbf{\tilde{H}}^{-1}\bm{\theta}_{k}$. Based on this transformation, we reparameterize the design matrix as
$\mathbf{\tilde{X}}\triangleq\left[\text{diag}(\mathbf{x}_{1})\mathbf{\tilde{H}}^{-1},...,\text{diag}(\mathbf{x}_{p})\mathbf{\tilde{H}}^{-1}\right].$ (10)
$L_{\text{FLAT}}(\bm{\theta})=\frac{1}{2}\|\mathbf{y}-\mathbf{\tilde{X}}\bm{\theta}\|_{2}^{2}+\lambda_{1}\sum_{k=1}^{p}\|\bm{\theta}_{k}\|_{1},$ (11)
where $\bm{\theta}\triangleq[\bm{\theta}_{1}^{\top},...,\bm{\theta}_{p}^{\top}]^{\top}\in\mathbb{R}^{np\times 1}$. We define the FLAT estimator of the coefficients as $\hat{\bm{\theta}}_{\mathrm{FLAT}}=\arg\min_{\bm{\theta}\in\Theta}L_{\text{FLAT}}(\bm{\theta})$. We can obtain the FLAT estimator of original regression coefficients using $\hat{\bm{\beta}}_{\mathrm{FLAT}}=\mathbf{\tilde{H}}^{-1}\hat{\bm{\theta}}_{\mathrm{FLAT}}$.
By minimizing (11), $\hat{\bm{\theta}}_{\mathrm{FLAT}}$ can detect the discrete spatial heterogeneity in effects while estimating the continuous spatial heterogeneity, and also select high-dimensional spatial features. To indicate non-convex spatial clusters in terms of feature effects, some classic density-based clustering algorithm (Ester et al., 1996; Miller and Han, 2009) can be used to further obtain the spatial surface in specific applications, such as the inverse temperature-salinity surface in thermohaline circulation. The details of the FLAT algorithms are summarized in Algorithm 1.
2:  Initialization of pairwise distances via (3) and obtain the adaptive weights $\pi_{(s_{i},s_{j})}$ in (2)
@@
3:  Obtain the fused structure $\bm{H}$ in (4) based on edge set $\mathbb{E}$, which is built by minimum spanning tree from Prim’s algorithm
6:  repeat
7:     for all  $j=1$ to $np$ do
8:        Update via soft-thresholding operator $\mathcal{S}_{\lambda}(x)=\text{sign}(x)(|x|-\lambda)_{+}$ and
$\theta_{j}^{(t+1)}=\mathcal{S}_{\lambda_{1}}\left(\theta_{j}^{(t)}-\eta\sum_{i=1}^{n}\tilde{X}_{ij}(y_{i}-\hat{y}_{i}^{(t)}+\tilde{X}_{ij}\theta_{j}^{(t)})\right)$
9:     end for
10:  until $\|\bm{\theta}^{(t+1)}-\bm{\theta}^{(t)}\|_{2}<10^{-6}$
13:  for all $(\epsilon,minPts)$ do
14:     Initialize all locations as unvisited
15:     for all unvisited locations $s_{i}$ do
17:        if $|N_{\epsilon}(s_{i})|\geq minPts$ then
18:           Create new cluster: $cluster.id\leftarrow cluster.id+1$
19:           for all $s_{j}\in N_{\epsilon}(s_{i})$ do
20:              Compute $N_{\epsilon}(s_{j})$
21:              if $|N_{\epsilon}(s_{j})|\geq minPts$ then
22:                 Merge $N_{\epsilon}(s_{j})$ into $N_{\epsilon}(s_{i})$
23:              else
24:                 Mark $s_{j}$ as noise
25:              end if
26:           end for
27:           Expand cluster: Add $N_{\epsilon}(s_{i})$ to cluster $C_{cluster.id}$
28:           Mark all locations in $C_{cluster.id}$ as visited
29:        end if
30:     end for
31:     Evaluate and select the clustering labels
32:  end for
Algorithm 1 FLAT Algorithm with Adaptive MST

### v2

@@ Abstract
This article introduces a new methodology model both discrete and continuous spatial heterogeneity simultaneously with an application in detection of hyper-plain in thermohaline circulation. To enable the data-driven detection of spatial boundaries with heterogeneity, we constructs an adaptive minimum spanning tree guided by both spatial proximity and coefficient dissimilarity, and combines both a spatial fused regularization and LASSO-type regularization to estimate the spatial coefficients under the framework of spatial regression. Numerical simulations demonstrate the effectiveness of proposed method in both estimation and heterogeneity detection. The usefulness of the approach is further illustrated via an analysis of oceanic data that provides new empirical finds about Atlantic with detected surfaces in temperature-salinity relationship.
@@
Fused regularization, Minimum spanning tree, Spatial heterogeneity, Temperature-salinity relationship, Thermohaline circulation
@@ 1 Introduction
Spatial heterogeneity commonly exist in multiple disinclines and plays a pivotal role in oceanic and earn system. The analyses of spatial heterogeneity provide critical information about the spatially varying relationships, such as soil moisture-temperature relationships across space (Zhang et al., 2024) and sudden shifts in rainfall regimes along mountain ridges (Luo et al., 2021). In ocean science, the temperature- salinity (T-S) relationship in thermohaline circulation are also widely studied in oceanic studies (Talley, 2011, 2013; Li and Yang, 2021).
One of the key challenges to model the spatial heterogeneity is to exhibits a dual nature of spatial non-stationarity of a complex system, with both smooth variations within specific regions and discrete transitions at their boundaries (Fotheringham et al., 2017). They are also known as continuous spatial heterogeneity (Bailey et al., 1995) and discrete heterogeneity (Ferrari et al., 2001; Anselin, 2010) respectively.
@@
To accurately model spatial heterogeneity, several models were developed (LeSage and Pace, 2009; Ward and Gleditsch, 2018) under the framework of spatial regression, where the local spatial dependencies are characterized by the a set of spatial coefficients (Brunsdon et al., 1996), such as the geographically weighted regression (GWR) model (Brunsdon et al., 1996), the spatial lag model (Lam and Souza, 2020), and the spatial Durbin model (Anselin, 2013). However, these models can only capture the continuous spatial heterogeneity without extra design to detect potential spatial boundaries, such as the inverse T-S surface in thermohaline circulation. Although there are empirical studies that specially designed two-step procedures to capture the discrete spatial heterogeneity (Chen et al., 2024; Duque et al., 2011) with clustering-based segmentation techniques (Raudenbush, 2002; Van Dusen and Nissen, 2019), they are not general data-driven methods for modeling spatial heterogeneity.
However, the detection of boundaries with sharp changes in spatial coefficients are important in many applications. As the T-S relationships is the primary driver of thermohaline circulation that plays a crucial role in regulating Earth’s climate, shaping marine ecosystems, and influencing sea-level rise (Rahmstorf, 2003; Velasco et al., 2021; Lin et al., 2023). Other examples of the discrete spatial heterogeneity can be widely find in studies of spatial modeling, such as economic boundaries (Gelfand et al., 2003) and areal wambling (Kudela et al., 2015; Luo et al., 2021). Real-world geographic spaces often exhibit both continuous and discrete spatial heterogeneity, but there are few general data-driven methods proposed for general modeling local smoothness while also identifying spatial boundaries adaptively.
To generally model the spatial heterogeneity with the dual nature, we propose a new approach, namely Fused Lasso with an Adaptive-minimum-spanning Tree (FLAT), which can effectively estimate the spatial coefficients in spatial regression model while capturing the discrete spatial heterogeneity by detecting the spatial boundaries simultaneously. Specifically, FLAT constructs fused penalization terms guided by minimum spanning tree structures that incorporate both spatial proximity and coefficient similarity, and introduces a structured adaptive regularization to model both continuous and discrete spatial heterogeneity in a unified optimization procedure. In addition, FLAT utilizes a density-based spatial clustering method to further detect the non-convex spatial regions with heterogeneity in applications. The comprehensive simulations demonstrate the dominant performance of FLAT in both parameter estimation and the detection of boundaries. A real-world oceanic application to model thermohaline circulation with new empirical finds demonstrates the effectiveness of FLAT in practice.
@@
The rest of the article is organized as follows. Section 2 introduces the spatial regression modal and the methodology of FLAT. Section 3 presents the simulation studies showing the effectiveness of FLAT. A real application of FLAT on thermohaline circulation is shown in Section 4 with analysis of new finds of North Atlantic Ocean. Section 5 concludes this article with a discussion. The technical details of both the algorithm and the real data analysis are given in Appendix.
@@ 2 Fused Lasso Regression With Adaptive Minimum Spanning Tree
In this paper, we focus on modeling the complex relationships and heterogeneity in spatial data $\{\bm{x}(s_{i}),y(s_{i})\}$ for $i\in[n]=\{1,\cdots,n\}$, where $\bm{x}(s_{i})=[x_{1}(s_{i}),\cdots,x_{p}(s_{i})]^{\top}$ is a $p$-dimensional vector of spatial features at the spatial location $s_{i}$, and $y(s_{i})$ is the corresponding response. Then we consider a spatial regression model (Ward and Gleditsch, 2018) with spatially varying coefficient as below:
$y(s_{i})=\bm{\beta}^{\top}(s_{i})\bm{x}(s_{i})+\epsilon(s_{i}),$ (1)
where $\bm{\beta}(s_{i})=[\beta_{1}(s_{i})\cdots,\beta_{p}(s_{i})]^{\top}$ is the heterogeneous coefficient vector at spatial location $s_{i}$, $\epsilon(s_{i})$ is independent random error term with mean 0 and variance $\sigma^{2}$. As introduced in Section 1, discrete spatial heterogeneity is widely considered in empirical analysis of spatial modeling, which is captured by the spatial clusters among $\bm{\beta}(s_{i})$ in spatial regression models.
To fully capture the heterogeneity in spatial data and detect the potential clusters with spatial boundaries, we consider to capture the spatial dependency between coefficients utilizing the minimum spanning tree, and then construct the fused regularization for estimation.
@@
To generate the minimum spanning tree $\mathbb{E}$ considering both spatial distance and the spatial heterogeneity, we firstly generated the spatial linkage structure $\mathbb{E}_{0}$ capturing the pure spatial distance spatial based on spatial distances $d(s_{i},s_{j})=\|s_{i}-s_{j}\|_{2}$. Then we obtain the estimator $\hat{\bm{\beta}}^{0}(s_{i}),i\in[n]$ for parameter $\beta(s_{i})$ at spatial location $s_{i}$ considering pure spatial distance by optimizing
@@
$\mathop{\arg\min}\limits_{\bm{\beta}}\frac{1}{n}\sum\limits_{i=1}^{n}\left(y(s_{i})-\bm{\beta}^{\top}(s_{i})\bm{x}(s_{i})\right)^{2}+\lambda\sum\limits_{k=1}^{p}\sum\limits_{(s_{i},s_{j})\in\mathbb{E}_{0}}|\beta_{k}(s_{i})-\beta_{k}(s_{j})|.$ (2)
To fully extract the information of spatial heterogeneity, we generate the full graph $\mathbb{E}_{1}$ among spatial locations based on the distance $\tilde{d}(s_{i},s_{j})$ defined as the $L_{2}$-norm of the difference in parameters between spatial locations, that is
$\tilde{d}(s_{i},s_{j})=\|\hat{\bm{\beta}}^{0}(s_{i})-\hat{\bm{\beta}}^{0}(s_{j})\|_{2}.$ (3)
@@
Then the minimum spanning tree $\mathbb{E}$ is obtained from $\mathbb{E}_{1}$ via Prim’s algorithm. Based on the definition of minimum spanning tree, the spatial linkage structure $\mathbb{E}$ will contain n1n-1 links.
To be consistent with the classic literature (Zhang et al., 2024; Li and Sang, 2019), we construct the connectivity matrix $\bm{H}\in\mathbb{R}^{(n-1)\times n}$ from $\mathbb{E}$, where element $\bm{H}_{l,i}$ of $\bm{H}$ on $l$-th row and $i$-th column os defined based on the connection status of vertices $i\in[n]$ along the edges $l\in[n-1]$. If the $l$-th potential edge, which starts from of spatial location $s_{i}$ to $s_{j}$, exists in the spatial linkage structure $\mathbb{E}$, then we set $\bm{H}_{l,i}=1$, $\bm{H}_{l,j}=-1$, and set all the other elements in $l$-th row of matrix $\bm{H}$ as 0. To capture discrete spatial heterogeneity in the spatial regression, we utilize the representation of minimum spanning tree above and consider the following regularization
$\|\bm{H}\bm{\beta}_{k}\|_{1}=\sum_{(s_{i},s_{j})\in\mathbb{E}}|\bm{\beta}_{k}(s_{i})-\bm{\beta}_{k}(s_{j})|,$ (4)
for $k$-th dimension of the spatial coefficient $\bm{\beta}$, where $\{\bm{\beta}_{k}=(\beta_{k}(s_{1}),...,\beta_{k}(s_{n})\}^{\top},\quad k=1,...,p$ represent the coefficients of the $k$-th explanatory variable among all spatial locations. Thus, $\|\bm{H}\bm{\beta}_{k}\|_{1}$ represents regularization based on the adaptive minimum spanning tree $\mathbb{E}$ in terms of the $k$-th spatial features.
To further includes the continuous spatial heterogeneity with the regularization defined in (4), we further denote $\pi_{(s_{i},s_{j})}=1/\tilde{d}(s_{i},s_{j})^{\gamma}$ as the adaptive weights between any two spatial $(s_{i},s_{j})$ for all $(s_{i},s_{j})\in\mathbb{E}$ and 0 otherwise. Although $\gamma$ can be any positive constant to control the adaptiveness of regularization, we follow the recommendation of Zou (2006) and set $\gamma=1$ in both simulations in Section 3 and real-data analysis in Section 4. Then we propose the objective function of FLAT as
$\hat{\bm{\beta}}_{\mathrm{FLAT}}=\arg\min_{\beta}\frac{1}{n}\sum_{i=1}^{n}\left\{y(s_{i})-\bm{\beta}^{\top}(s_{i})\bm{x}(s_{i})\right\}^{2}+\lambda_{1}\sum_{k=1}^{p}\|\bm{\pi}\bm{H}\bm{\beta}_{k}\|_{1}+\lambda_{2}\sum_{k=1}^{p}\|\bm{\beta}_{k}\|_{1},$ (5)
where $\lambda_{1}$ and $\lambda_{2}$ are tuning parameters controlling the strength of penalization to capture the discrete spatial heterogeneity and the sparsity in spatial features. The adaptive weights $\bm{\pi}$ further enhance the detection of spatial heterogeneity while reducing the estimation error for the spatial locations with homogeneity. With the specially designed regularization based on the adaptive-minimum-spanning tree, we can effectively detect the discrete spatial heterogeneity and select the spatial features simultaneously while estimate the continuous spatial coefficients by optimizing (5). The details of the derivation and the algorithms are given in Section 5.1 of Appendix, and the R tools for FLAT are publicly available [^1] https://github.com/ZhaoyuXingStat/FLAT. We next investigate the numerical performance of $\hat{\bm{\beta}}_{\mathrm{FLAT}}$ in the following sections.